OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values

jimthompson5910 (jim_thompson5910):

Plug in x = 0 y = -2x - 3 y = -2(0) - 3 y = 0 - 3 y = -3 So when x = 0, y is y = -3

jimthompson5910 (jim_thompson5910):

Now plug in x = 1 y = -2x - 3 y = -2(1) - 3 y = -2 - 3 y = -5 So when x = 1, y is y = -5

jimthompson5910 (jim_thompson5910):

Plug in x = 2 y = -2x - 3 y = -2(2) - 3 y = -4 - 3 y = -7 So when x = 2, y is y = -7 Etc etc. You can keep going to make the table as big as you want. You only need 2 points, but 3 is also a good number to have This gives us the table x | y -------- 0 | -3 1 | -5 2 | -7

OpenStudy (anonymous):

So, all of this is the answer @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it's the steps and the answer

jimthompson5910 (jim_thompson5910):

The table is the final answer, but everything else is a lead up to it

jimthompson5910 (jim_thompson5910):

if you need a bigger table, then you would get x | y -------- 0 | -3 1 | -5 2 | -7 3 | -9 4 | -11 5 | -13

OpenStudy (anonymous):

A bigger tables isnt necessary is it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

it depends on what your teacher/book is looking for, but technically, you can get away with having 2 rows since you need at least 2 points

jimthompson5910 (jim_thompson5910):

I would do 3 just because I've seen some teachers use 3

OpenStudy (anonymous):

Oh ok @jim_thompson5910, Thank you so much

jimthompson5910 (jim_thompson5910):

you're welcome